OpenStudy (anonymous):

points (4,2) and (5,1) lie on line n. what is the slope of the line that is parallel to j? A.0 B.1 C.-1 D.Undefined

OpenStudy (saifoo.khan):

where's j

OpenStudy (anonymous):

points (4,2) and (5,1) lie on line j. what is the slope of the line that is parallel to j? A.0 B.1 C.-1 D.Undefined

OpenStudy (saifoo.khan):

if you find the slope, \[slope = \frac{y_2 - y_1}{x_2-x_1} \to \frac{1-2}{5-4}\]

OpenStudy (saifoo.khan):

Can you solve?

OpenStudy (anonymous):

na man not good at all with these type of crap only have 2 questions left

OpenStudy (saifoo.khan):

it's simple, 1-2 = -1 same as with denominator.

OpenStudy (anonymous):

okay give me the answers for theses last two please ill be sooo grateful points (1,1) and (1,4) lie on line k. what is the slope of the line that is parallel to k? A.0 B.1 C.-1 D.Undefined

OpenStudy (saifoo.khan):

Direct answers is marked as cheating. sorry.

OpenStudy (anonymous):

points (0,0) and (3,3) lie on line k. what is the slope of the line that is perpendicular to k? A.1 B.-1 C.0 D.Undefined